OpenStudy (anonymous):

What is the value of x? 4(6 + 7) = 4 • 6 + x • 7 need help

OpenStudy (anonymous):

x=4

OpenStudy (anonymous):

x = 4 4(6+7) = 4 x 6 + x x 7 52 = 24 + 7x 28 = 7x x = 4

OpenStudy (anonymous):

so if you distribute 4(6+7) you would get 24+28 6*4=24 and x*7=28 28/7= 4

OpenStudy (anonymous):

thank you soo much guys math isnt really my subject im very bad in math

OpenStudy (anonymous):

No problem

OpenStudy (anonymous):

any time!!!